What is the probability of obtaining a six at least once in two rolls of a die?
1
Expert's answer
2019-11-04T09:28:54-0500

P(obtaining 6 at least once) = 1 - P(no obtaining 6)


P(no obtaining 6 in two rolls) = 5 / 6 * 5 /6 = 25 / 36


P(obtaining 6 at least once) = 1 - 25 / 36 = 11/36


Answer: 11/36